An annular ring with inner and outer R1 and R2 is rolling without slipping with a uniform angular speed The ratio of the force experienced by the particles situated on inner and outer parts of the ring F1/F2 is

Grade:12

3 Answers

Asit Avilash
7 years ago
As the ring is rolling without slipping,Then V=w^2*r( w- angular velocity and r- radius). Hence, inside particle will have a velocity v1=w^2*r1 and outside particle will have v2=w^2*r2We know force,F=mv^2/rNow F1/F2=(m*v1^2/r1)/ (m*v2^2/r2)Substituting the value of v1 and v2 from above, we get F1/F2= r1/r2
Asit Avilash
7 years ago
As the ring is rolling without slipping,Then V=w*r( w- angular velocity and r- radius). Hence, inside particle will have a velocity v1=w*r1 and outside particle will have v2=w*r2We know force,F=mv^2/rNow F1/F2=(m*v1^2/r1)/ (m*v2^2/r2)Substituting the value of v1 and v2 from above, we get F1/F2= r1/r2
Asit Avilash
7 years ago
As the ring is rolling without slipping,Then V=w*r( w- angular velocity and r- radius). Hence, inside particle will have a velocity v1=w*r1 and outside particle will have v2=w*r2We know force,F=mv^2/r. Now F1/F2=(m*v1^2/r1)/ (m*v2^2/r2). Substituting the value of v1 and v2 from above, we get (F1/F2= r1/r2) ans.